# PACOX: A FPGA-based Pauli Composer Accelerator for Pauli String Computation

### v1

@@ IV Experiments and Results
We further compare the proposed PACOX with widely used CPU-based quantum software libraries on standard testcases, including Qiskit [16] and PennyLane [4]. Figure 6 shows the execution time for computing $n$-qubit Pauli strings on the Two-local[23], VQE [26], TFIM [28], and Stabilizer code [13] datasets with qubit counts ranging from 3 to 19. Across all test cases, PACOX consistently outperforms Qiskit and PennyLane, and the performance gap widens as the $n$ increases. This highlights the effectiveness of the proposed hardware architecture for accelerating practical Pauli-based workloads. Figure 6 also reports the measured PDP for PACOX and the CPU-based libraries over the same qubit range. The PACOX achieves significantly lower PDP values across all datasets, demonstrating that its performance gains are obtained with superior energy efficiency.
@@ V Conclusion
This paper presents PACOX, a dedicated FPGA-based architecture for accelerating Pauli string computation in hybrid quantum-classical systems. PACOX employs a compact binary encoding that transforms the entire process into an XOR-based index permutation and phase accumulation, enabling efficient parallel and pipelined execution on an FPGA. Implemented on a Xilinx ZCU102 platform, PACOX achieves up to two orders of magnitude speedup over representative CPU-based Pauli computation methods, while significantly reducing memory usage and power consumption. These results demonstrate that PACOX provides a high-speed and energy-efficient hardware solution for accelerating Pauli-based quantum workloads.

### v2

@@ IV-D Comparison with CPU-Based Software
We further compare the proposed PACOX with widely used CPU-based quantum software libraries on standard datasets, including Qiskit [9] and PennyLane [3] conducted on an Intel 3.30 GHz i9-10940X processor. Fig.  7 shows the execution time for computing $n$-Pauli strings on the Heisenberg [12], QChem [13], TFIM [15], and Stabilizer [5] testcases with qubit counts ranging from 3 to 19. All results correspond to real executions on both platforms. Across all datasets, PACOX consistently outperforms Qiskit and PennyLane, and the performance gap widens as the number of qubits increases. This highlights the effectiveness of the proposed hardware architecture for accelerating practical Pauli-based workloads. Fig. 7 also reports the measured PDP for PACOX and the CPU-based libraries over the same qubit range. PACOX achieves significantly lower PDP values across all datasets, demonstrating that its performance gains are obtained with superior energy efficiency.
Fig. 8 shows the total runtime of the PACOX system for varying values of $n$ from 3 to 19. As the $n$ increases, both execution time and total runtime (including data transfer time) increase. However, the data transfer time grows much more significantly than the execution time. This happens because PACOX uses 32 PEs to compute 32 input data points simultaneously. But, the ZCU102 FPGA’s DMA can only transfer 4 output data points at a time from the PL to the PS. Specifically, the bandwidth of the ZCU102 FPGA is only 4 GB/s working at a frequency of 250 MHz. Additional time is also added due to tasks such as processing communications and sending initial signals from the PS to the PL. As a result, the total runtime increases much more than the execution time. This bottleneck in data transfer could be alleviated by using the Alveo U280 FPGA, which offers high-bandwidth memory of up to 64 GB/s at a frequency of 250 MHz.
@@ V Conclusion
This paper presents PACOX, a dedicated FPGA-based architecture for accelerating Pauli string computation in hybrid quantum-classical systems. PACOX employs a compact binary encoding that transforms the entire process into an XOR-based index permutation and phase accumulation, enabling efficient parallel and pipelined execution on an FPGA. Implemented on a Xilinx ZCU102 platform, PACOX outperform representative CPU-based Pauli computation methods, while significantly reducing memory usage and power consumption. These results demonstrate that PACOX provides a high-speed and energy-efficient hardware solution for accelerating Pauli-based quantum workloads.
Acknowledgment
This work was supported by JST-ALCA-Next (JPMJAN23F4), the NAIST Senju Monju Project (Daiichi-Sankyo ”Habataku” Support Program for the Next Generation of Researchers).